OpenStudy (anonymous):

Help ASAP please:) Will fan & Medal!! Thank you:) Match each set with its description. (Points : 1) Potential Matches: 1 : {0, 1, 2, 3, ...} 2 : {...-3, -2, -1, 0, 1, 2, 3,...} 3 : A number which can be written as a/b. 4 : {1, 2, 3,...} Answer : Rational Numbers : Whole Numbers : Integers : Natural Numbers

OpenStudy (anonymous):

@marsss Thank you!! :) I have a few other questions if you could help

OpenStudy (anonymous):

Mark every set that applies in the remaining questions. Classify -5. (Points : 1) Whole Number Integer Rational Number Irrational Number

OpenStudy (anonymous):

yes but i think my answer was wrong lol

OpenStudy (anonymous):

Lol do you know which ones were wwrong?

OpenStudy (anonymous):

you should exchange whole numbers and natural numbers

OpenStudy (anonymous):

and -5 is an integer

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

-5 is also a rational number

OpenStudy (anonymous):

Classify 8. (Points : 1) Whole Number Integer Rational Number Irrational Number Question 7. 7. Classify -3.4 (Points : 1) Whole Number Integer Rational Number Irrational Number

OpenStudy (anonymous):

8 is a whole number, an integer and a rational number -3.4 is a rational number

OpenStudy (anonymous):

okay thank you:) Let me get the other questuions

OpenStudy (anonymous):

Classify 0.626539212... (Points : 1) Whole Number Integer Rational Number Irrational Number Question 10. 10. Classify 0.767676... (Points : 1) Whole Number Integer Rational Number Irrational Number

OpenStudy (anonymous):

@marsss

OpenStudy (anonymous):

0.626539212... is a irrational number 0.767676... is a rational number because it is 76/99

OpenStudy (anonymous):

Thank you:) Here is the last one Five is a rational number. (Points : 1) True False

OpenStudy (anonymous):

of course true :D

OpenStudy (anonymous):

Thank you so so much!!

OpenStudy (anonymous):

you're welcome :D

OpenStudy (anonymous):

If your not busy would you be able to help me with a few more? @marsss

OpenStudy (anonymous):

yes, if i can

OpenStudy (anonymous):

okay thank you!!

OpenStudy (anonymous):

-100 (50) = (Points : 1) 5,000 -5,000 150 -2 Question 2. 2. -12 (-3) = (Points : 1) 36 -36 -15 4 Question 3. 3. 16 ÷ (-2) = (Points : 1) -32 8 14 -8 Question 4. 4. -30 ÷ (-10) (Points : 1) 300 3 -20 -3

OpenStudy (anonymous):

3. 16 ÷ (-2) = (Points : 1) -32 8 14 -8 Question 4. 4. -30 ÷ (-10) (Points : 1) 300 3 -20 -3

OpenStudy (anonymous):

first answer is -5000 second answer is 36 but what does this ������������������ mean?

OpenStudy (anonymous):

I have no idea. let me post it again and see what happend. Lol

OpenStudy (anonymous):

3. 16 ÷ (-2) = (Points : 1) -32 8 14 -8 Question 4. 4. -30 ÷ (-10) (Points : 1) 300 3 -20 -3

OpenStudy (anonymous):

Ughh

OpenStudy (anonymous):

does it mean dividing?

OpenStudy (anonymous):

16 ÷ (-2) = -32 8 14 -8

OpenStudy (anonymous):

16 divided by (-2)

OpenStudy (anonymous):

so question 3 is -8 question 4 is 3

OpenStudy (anonymous):

Carmella is an accountant for a business. She is making a report to investors. Last year the business lost $32,400. What was the average loss per month? (Points : 1) -$2,700 -$3,240 -$324 -$270

OpenStudy (anonymous):

you must divide -32400 by 12 months -32400/12= -2700

OpenStudy (anonymous):

Thank you!! Evaluate 3wz if w = -2 and z = 4. (Points : 1) 5 -24 -14 10 Question 10. 10. 24 - 2 x 8 = (Points : 1) 8 176 -96

OpenStudy (anonymous):

3wz means 3*w*z=3*(-2)*4=-24 you should do the multipication first 24-2*8=24-16=8

OpenStudy (anonymous):

Thank you :) What about this one? Evaluate 4x - 3y if x = -3 and y = 4. (Points : 1) -8 8 24 -24

OpenStudy (anonymous):

4*x-3*y=4*(-3)-3*4=-12-12=-24 how can't you solve those? how old are you?

OpenStudy (anonymous):

12.. Were just learning these and I'm going through youtube and other resources to figure these out

OpenStudy (anonymous):

lol ok :D

OpenStudy (anonymous):

Simplify: -(3x - 2). (Points : 1) 2 - 3x -3x - 2 3x - 2 2 + 3x Question 18. 18. Simplify: -(x + 8). (Points : 1) x + 8 -x - 8 x - 8 8 - x

OpenStudy (anonymous):

-(3x-2)=-3x+2=2-3x -(x+8)=-x-8

OpenStudy (anonymous):

 25 10 -25 -10

OpenStudy (anonymous):

is there any brackets?

OpenStudy (anonymous):

Thank you so much for helping:)

OpenStudy (anonymous):

no brackets

OpenStudy (anonymous):

so the answer is -25. but if -5 would be in paranthesis, the answer would be 25

OpenStudy (anonymous):

 10 1,450 -40 2,500 No brackets